OpenStudy (anonymous):

Price: $39,145 Current interest rate: 3% Using the function A(t)=P(1+r/n)^nt, create the function that represents your new car loan that is compounded monthly. The principle will be the price of the vehicle you selected, not how much you are putting down. Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal.

OpenStudy (anonymous):

Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P[(1+r/n)1c]^cnt where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1+r/n.

OpenStudy (anonymous):

i have done number 1 and 2. need help on three.

OpenStudy (anonymous):

My answers: 1.A(t)=39,145(1+0.03/12)^12t 2.-45.5

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

please, can someone help me!!! :(

OpenStudy (anonymous):

any idea @wio?

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

PLEASE HELP ME!!!! IS THERE SOMEBODY OUT THERE THAT CAN HELP ME!!!!! REALLY REALLY NEED IT!!!!!!! :(

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

I have an idea of how to do it.

OpenStudy (anonymous):

Though I don't understand your formula.

OpenStudy (anonymous):

whats wrong with it?

OpenStudy (anonymous):

I've never seen that formula before, and don't know where it comes from.

OpenStudy (anonymous):

This site seems to talk about a similar problem: https://prezi.com/olovfcq7zpwg/704-solving-exponential-equations-with-unequal-bases/

OpenStudy (anonymous):

yea that were i get stuck at. i saw this before, and i really didn't understand that ether

OpenStudy (anonymous):

this is how i plugged in my problem:⦁ A(t)=39,145[(1+0.03/12^1/4)^4*12*t

OpenStudy (anonymous):

i just don't know where to go from here.

OpenStudy (anonymous):

Here is how you would solve for \(t\)\[ A(t)=P(1+r/n)^{nt} \\ A/P=(1+r/n)^{nt} \\ \ln(A/P) = nt\ln(1+r/n)\\ t = \frac{\ln(A/P)}{n\ln(1+r/n)} \]

OpenStudy (anonymous):

They give you \(A=10,000.00\)

OpenStudy (anonymous):

And you know the rest of the variables, so plug them in to get \(t\).

OpenStudy (anonymous):

i don' know what to do with the 1/4th.

OpenStudy (anonymous):

Here is how you would solve for \(t\)\[ A(t)=P\left(\frac1c+\frac r{cn}\right)^{cnt} \\ A/P=\left(\frac1c+\frac r{cn}\right)^{cnt} \\ \ln(A/P) = cnt\ln\left(\frac1c+\frac r{cn}\right)\\ t = \frac{\ln(A/P)}{cn\ln\left(\frac1c+\frac r{cn}\right)} \]

OpenStudy (anonymous):

thanks!!!!! :)

OpenStudy (anonymous):

@wio how did u put that in that format?

OpenStudy (anonymous):

i would love to know how to do it :)